# Low-dimensional tori in Calogero–Moser–Sutherland systems

### v1

@@ 4. Action-angle variables
The cotangent bundle $T^{*}G$ carries the natural symplectic form
@@
Here we used the trivialization $T^{*}G\simeq\mathfrak{g}^{*}\times G$ by right translation. The identification $\mathfrak{g}^{*}\simeq\mathfrak{g}$ is given by the Killing form. We also assume that $G$ is a compact matrix group.
This symplectic form is $G$-invariant with respect to the lift of the conjugation action to $T^{*}G$. After the trivialization and identification $\mathfrak{g}^{*}\simeq\mathfrak{g}$, it acts as
$h\colon\quad x\mapsto hxh^{-1},\quad g\mapsto hgh^{-1}.$
The symplectic form (62) restricted to $\mu^{-1}(\mathcal{O})$ gives a presymplectic form, which remains $G$-invariant, and becomes a symplectic form $\Omega_{\mathcal{O}}$ on $S(\mathcal{O})=\mu^{-1}(\mathcal{O})/G$. We can model $S(\mathcal{O})$ as the space of pairs $(x,g)$, up to the action $H\subset SU(n)$, where $x=(x_{1}\geq x_{2}\geq\ldots\geq x_{n})$, $\sum_{i=1}^{n}x_{i}=0$, is diagonalized. On the big stratum $g=\alpha X$ (33), on others strata $g=\alpha PY$, see remark 3.
Now, let us describe the restriction $\Omega_{\mathcal{O}}$ of the symplectic form (62) in terms of $(y,\theta)$ coordinates.
@@
On the generic stratum
(63) $\Omega_{\mathcal{O}}=-\sum_{j=1}^{n}\mathrm{d}x_{j}\wedge\mathrm{d}\eta_{j}.$
@@
Using the representation (33) for $\Omega$ form we get
@@
(65) $\mathrm{Tr}\left(x\,\mathrm{d}X\,X^{-1}\right)=\sum_{i,j=1}^{n}x_{i}\,\mathrm{d}X_{ij}\cdot(X^{t})_{ji}=\frac{1}{2}\sum_{i,j=1}^{n}x_{i}\,\mathrm{d}(X_{ij})^{2}=\frac{1}{2}\sum_{i=1}^{n}x_{i}\,\mathrm{d}\sum_{j=1}^{n}(X_{ij})^{2}=0,$
where we take into account the orthogonality of $X$, which implies $\sum_{j=1}^{n}(X_{ij})^{2}=1$.
Thus
@@
Taking into account constraints $\sum_{j=1}^{n}\mathrm{d}x_{j}=0$ and $\sum_{j=1}^{n}\mathrm{d}\eta_{j}=0$ we can write $\Omega_{\mathcal{O}}$ as
(67) $\Omega_{\mathcal{O}}=-\sum_{j=2}^{n}\mathrm{d}y_{j}\wedge\mathrm{d}\theta_{j}.$
@@
On a $2s$-dimensional stratum of $S(\mathcal{O})$ the symplectic form $\Omega_{\mathcal{O}}$ becomes
(68) $\Omega_{\mathcal{O}}=-\sum_{a=1}^{s}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\theta_{j_{a}}.$
@@
Using representation (47) for $\Omega$ form we get
@@
Define $\alpha_{j_{a}}=e^{i\eta_{j_{a}}}$ for $a=0,\ldots,s$. Constrains $\prod_{a=0}^{s}\alpha_{j_{a}}=1$ implies.
@@
(71) $\Omega_{\mathcal{O}}=i\sum_{a=0}^{s}\mathrm{d}x^{P}_{j_{a}}\wedge\alpha_{j_{a}}^{-1}\mathrm{d}\alpha_{j_{a}}=\\ -\sum_{a=1}^{s}\frac{1-j_{a}}{n}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\eta_{1}-\sum_{a=1}^{s}\sum_{b=1}^{s}\left(\frac{n+1-j_{b}}{n}-\delta_{j_{b}-1\geq j_{a}}\right)\mathrm{d}y_{j_{b}}\wedge\mathrm{d}\eta_{j_{a}}=\\ -\sum_{a=1}^{s}\sum_{b=1}^{s}\left(1-\delta_{j_{b}-1\geq j_{a}}\right)\mathrm{d}y_{j_{b}}\wedge\mathrm{d}\eta_{j_{a}}=-\sum_{b=1}^{s}\mathrm{d}y_{j_{b}}\wedge\sum_{a=b}^{s}\mathrm{d}\eta_{j_{a}}.$
@@
Remark 5.
Note that $\theta$’s defined in proposition 3 differ slightly from $\theta$’s defined in proposition 2. This difference arises due to the introduction of a shift matrix $P$ (40).
The coordinates $(y_{j_{a}},\theta_{j_{a}})$ are Darboux action-angle coordinates on this $s$-dimensional stratum.
5. The multitime dynamics of Calogero–Moser–Sutherland system
5.1. The multi-time dynamics on $T^{*}G$.
Let $v_{k}$ be the Hamiltonian vector field on $T^{*}G\simeq\mathfrak{g}^{*}\times G$ generated by $H_{k}=\frac{1}{k}\operatorname{tr}(x^{k})$, i.e. $i_{v_{k}}\Omega=-dH_{k}$, where $\Omega$ is the symplectic form (62) on $T^{*}G$. We have
(72) $v_{k}=-i\sum_{i,j=1}^{n}\left(x^{k-1}g\right)_{ij}\frac{\partial}{\partial g_{ij}},$
Because Hamiltonians $H_{k}$ Poisson commute on $T^{*}G$, vector fields $v_{k}$ also commute. Their multi-time flows are solutions to the system of differential equations:
(73) $\frac{\mathrm{d}x({\bf t})}{\mathrm{d}t_{k}}$ $=0,$
(74) $\frac{\mathrm{d}g({\bf t})}{\mathrm{d}t_{k}}$ $=ix({\bf t})^{k-1}g({\bf t}).$
The initial value problem $x(0)=x,\,g(0)=g$ for these equations can be solved explicitly:
(75) $x({\bf t})=x,\qquad g({\bf t})=e^{i\sum_{k=1}^{n}t_{k}\,x^{k-1}}g.$
This dynamics preserves Hamiltonians $H_{k}$ and the moment map $\mu(x({\bf t}),g({\bf t}))=\mu(x,g)=x-g^{-1}xg$.
5.2. The multi-time dynamics on the generic stratum.
In coordinates $(y,\theta)$ the symplectic form on $S(\mathcal{O})$ obtained by the Hamiltonian reduction from (62) is
which means Poisson brackets between coordinate function $\{y_{i}\}$ and $\{\theta_{i}\}$ are
$\{y_{i},y_{j}\}=0,\qquad\{\theta_{i},\theta_{j}\}=0,\qquad\{\theta_{i},y_{j}\}=\delta_{ij}.$
Any $G$-invariant function $H$ on $T^{*}G$ defines a $G$-invariant function on $\mu^{-1}(\mathcal{O})\subset T^{G}$ and a function $S(\mathcal{O})=\mu^{-1}(\mathcal{O})/G$ which we denote by same letter $H$. Hamiltonian vector fields $v_{k}$ restrict to $S(\mathcal{O})$ and have the form
(77) $v_{k}=\sum_{i,j}x_{i}^{k-1}\frac{\partial x_{i}}{\partial y_{j}}\frac{\partial}{\partial\theta_{j}},\qquad\text{ for }k=2,\ldots,n.$
Here we can express $x_{i}$ in terms of $y_{i}$ according to (36):
(78) $v_{k}=\sum_{j=2}^{n}u^{k}_{j}\frac{\partial}{\partial\theta_{j}},$
@@ 5. The multitime dynamics of Calogero–Moser–Sutherland system
(79) $u^{k}_{j}=\sum_{i=1}^{n}\left(\frac{1}{n}\sum_{j=2}^{n}(n+1-j)y_{j}-\sum_{j=2}^{i}y_{j}\right)^{k-1}\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right).$
Flow lines of these vector fields are linear in coordinates $\theta_{l}$. The multitime flow line with $y_{j}(0)=y_{j}$, $\theta_{j}(0)=\theta_{j}$ is
(80) $y_{j}({\bf t})=y_{j},\qquad\theta_{j}({\bf t})=\theta_{j}-\sum_{k=1}^{n-1}u^{k}_{j}\,t_{k},\qquad\text{ for }k=2,\ldots,n.$
From this we have:
(81) $x({\bf t})=x,\qquad g({\bf t})=\alpha(\mathbf{t})Y,$
where $\alpha_{k}(\mathbf{t})=\exp\left(i(\theta_{k+1}(\mathbf{t})-\theta_{k}(\mathbf{t}))\right)$ with $\theta_{1}(\mathbf{t})=\theta_{n+1}(\mathbf{t})=0$ and $Y$ is defined in (33).
The time dependence of coordinates $q_{j}$ is determined by $g(\mathbf{t})$ since $\{e^{iq_{j}(\mathbf{t})}\}$ are eigenvalues of $g(\mathbf{t})$.
5.3. The multi-time dynamics on low-dimensional strata.
When $\psi_{i}=0$ except for $i=j_{0},j_{1},\ldots,j_{s}$ (here $j_{0}=1$), the matrix $g$ can be written as
(82) $g=\alpha PY.$
In coordinates $(y,\theta)$ the symplectic form on the correspond stratum of $S(\mathcal{O})$, see remark 3, is
(83) $\Omega=-\sum_{a=1}^{s}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\theta_{j_{a}},$
which implies
$\{y_{j_{a}},y_{j_{b}}\}=0,\qquad\{\theta_{j_{a}},\theta_{j_{b}}\}=0,\qquad\{\theta_{j_{a}},y_{j_{b}}\}=\delta_{ab}.$
Hamiltonian vector fields $v_{k}$ restrict to $S(\mathcal{O})$ and have the form
(84) $v_{k}=\sum_{i=1}^{n}\sum_{a=1}^{s}x_{i}^{k-1}\frac{\partial x_{i}}{\partial y_{j_{a}}}\frac{\partial}{\partial\theta_{j_{a}}},\qquad\text{ for }k=2,\ldots,n.$
Here we can choose $s$ independent vector fields. In particular, we can choose the first $s$ vectors.
Here we can express $x_{i}$ in terms of $y_{i}$ according to (36):
(85) $v_{k}=\sum_{a=2}^{s}u^{k}_{a}\frac{\partial}{\partial\theta_{j_{a}}},$
where
(86) $u^{k}_{a}=\sum_{i=1}^{n}\left(\frac{n-1}{2}c-(i-1)c+\sum_{a=1}^{s}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right)(y_{j_{a}}-c)\right)^{k-1}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right).$
The solution to the initial value problem $y_{j_{a}}(0)=y_{j_{a}},\,\theta_{j_{a}}(0)=\theta_{j_{a}}$ for these equations is
(87) $y_{j_{a}}({\bf t})=y_{j_{a}},\qquad\theta_{j_{a}}({\bf t})=\theta_{j_{a}}-\sum_{k=1}^{s}u^{k}_{a}\,t_{k},\qquad\text{ for }a=1,\ldots,s.$
So,
(88) $x({\bf t})=x,\qquad g({\bf t})=\alpha(\mathbf{t})P\,Y,$
where $\alpha_{j_{a}}(\mathbf{t})=\exp\left(i(\theta_{j_{a}}(\mathbf{t})-\theta_{j_{a+1}}(\mathbf{t}))\right)$ with $\theta_{1}(\mathbf{t})=\theta_{j_{s+1}}(\mathbf{t})=0$, $Y$ is defined in (48) and $\alpha$ is defined in (50).
The time dependence of coordinates $q_{j}$ is determined by $g(\mathbf{t})$ since $\{e^{iq_{j}(\mathbf{t})}\}$ are eigenvalues of $g(\mathbf{t})$.
Example 3.
Zero-dimensional stratum.
On this stratum, particles are in equilibrium. Their position are terminated by eigenvalues of (53)
(89) $\mathrm{det}\left(\lambda-g\right)=\lambda^{n}+(-1)^{n}=0.$
Example 4.
One-dimensional strata.
There are $(n-1)$ such strata, labeled by $j=2,\ldots,n$, with coordinates $(y_{j},\theta_{j})$ defined in example 2.
Positions of particles are eigenvalues of $g(t)$ and thus are roots of the characteristic equation
(90) $\mathrm{det}\left(\lambda-g\right)=\lambda^{n}+(-1)^{j-1}q\beta^{-1}\lambda^{n-j+1}+(-1)^{n-j+1}q\beta\lambda^{j-1}+(-1)^{n}=0,$
where $\beta=\beta(t)=\alpha_{1}^{-1}=\alpha_{j}=\exp\left(i\theta_{j}(t)\right)=\exp(i\,(\theta_{j}-u\,t))$ and $10<q<$ 1 defined in example 2, and
(91) $u=\sum_{i=1}^{n}\left(\frac{n-1}{2}c-(i-1)c+\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)(y_{j}-c)\right)\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)=\\ (n-j+1)(j-1)\left(\frac{c}{2}+\frac{y-c}{n}\right).$
Appendix A Lemma about determinant of sum of matrix plus rank $1$ matrix
@@
Let $D$ be a diagonal matrix $D_{ab}=D_{a}\delta_{ab}$ and $P_{ab}=\varphi_{a}\psi_{b}$ be a $1$-dimensional projector. Then
(92) $\det(D-z-P)=\prod_{a}(D_{a}-z)-\sum_{\beta}\varphi_{b}\psi_{b}\prod_{\begin{subarray}{c}a\neq b\end{subarray}}(D_{a}-z)$
@@
(95) $=\prod_{\alpha}(D_{\alpha}-z)\exp\left(\sum_{k\geq 1}\frac{1}{k}\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}\right)$
@@
(97) $=\sum_{a_{1},\ldots,a_{k}}(D_{a_{1}}-z)^{-1}\cdots(D_{a_{k}}-z)^{-1}\prod_{i=1}^{k}\varphi_{a_{i}}\psi_{a_{i}}$
@@
$\exp\left(\sum_{k=1}^{\infty}\frac{1}{k}\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}\right)=1-\sum_{b=1}^{n}(D_{b}-z)^{-1}\varphi_{b}\psi_{b}$

### v2

@@ 4. Action-angle variables
Here we show that natural coordinates $(y,\theta)$ are Darboux action-angle variables on each $S(\mathcal{O})_{\{j\}}$.
The cotangent bundle $T^{*}G$ carries the natural symplectic form, which, after the trivialization of $T^{*}G$ by right translations can be written in terms of $(x,g)\in su(n)^{*}\times SU(n)\simeq T^{*}SU(n)$ as
@@
Here $x$ is an $n\times n$ Hermitian traceless matrix an $g$ is and unitary unimodular $n\times n$ matrix.
This symplectic form is $G$-invariant with respect to the lift of the conjugation action to $T^{*}G$. It acts on $(x,g)$ as
(85) $h\colon\quad x\mapsto hxh^{-1},\quad g\mapsto hgh^{-1}.$
The restriction of the symplectic form (84) to $\mu^{-1}(\mathcal{O})$ defines a presymplectic form, which remains $G$-invariant, and becomes a symplectic form $\Omega_{\mathcal{O}}$ on $S(\mathcal{O})=\mu^{-1}(\mathcal{O})/G$. We can model $S(\mathcal{O})$ as the space of pairs $(x,g)$, up to the action $H\subset SU(n)$, where $x=(x_{1}\geq x_{2}\geq\ldots\geq x_{n})$, $\sum_{i=1}^{n}x_{i}=0$, is diagonalized. On the maximal stratum $g=\alpha X$ (46), on other strata $g=\alpha PY$, see remark 2.
Now, let us describe the symplectic form $\Omega_{\mathcal{O}}$ in terms of $(y,\theta)$ coordinates introduced in section 3.4.1.
@@
On the maximal stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}$ the symplectic form $\Omega_{\mathcal{O}}$ is
(86) $\Omega_{\mathcal{O}}=-\sum_{j=1}^{n}\mathrm{d}x_{j}\wedge\mathrm{d}\eta_{j},$
@@
Evaluating the symplectic form $\Omega$ on the representative $(x,g=\alpha X)$ we obtain
@@
(88) $\mathrm{Tr}\left(x\,\mathrm{d}X\,X^{-1}\right)=\sum_{i,j=1}^{n}x_{i}\,\mathrm{d}X_{ij}\cdot(X^{t})_{ji}=\frac{1}{2}\sum_{i,j=1}^{n}x_{i}\,\mathrm{d}(X_{ij})^{2}=\frac{1}{2}\sum_{i=1}^{n}x_{i}\,\mathrm{d}\sum_{j=1}^{n}(X_{ij})^{2}=0.$
Here we took into account the orthogonality of $X$, which implies $\sum_{j=1}^{n}(X_{ij})^{2}=1$. Thus, on the maximal stratum
@@
After changing variables from $(x,\eta)$ to $(y,\theta)$ we can write $\Omega_{\mathcal{O}}$ as
@@
On the 2s2s–dimensional stratum $S(\mathcal{O})_{\{j\}}$ we have
(91) $\Omega_{\mathcal{O}}=-\sum_{a=1}^{s}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\theta_{j_{a}},$
@@
Writing $g=\alpha PY$ as in (63) for the restriction of $\Omega_{\mathcal{O}}$ we get
@@
Define $\alpha_{j_{a}}=e^{i\eta_{j_{a}}}$ for $a=0,\ldots,s$. Constrain $\prod_{a=0}^{s}\alpha_{j_{a}}=1$ implies
@@
(94) $\Omega_{\mathcal{O}}=i\sum_{a=0}^{s}\mathrm{d}x^{P}_{j_{a}}\wedge\alpha_{j_{a}}^{-1}\mathrm{d}\alpha_{j_{a}}=$- $\sum_{a=1}^{s}\frac{1-j_{a}}{n}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\eta_{1}-\sum_{a=1}^{s}\sum_{b=1}^{s}\left(\frac{n+1-j_{b}}{n}-\delta_{j_{b}-1\geq j_{a}}\right)\mathrm{d}y_{j_{b}}\wedge\mathrm{d}\eta_{j_{a}}=$
$-\sum_{a=1}^{s}\sum_{b=1}^{s}\left(1-\delta_{j_{b}-1\geq j_{a}}\right)\mathrm{d}y_{j_{b}}\wedge\mathrm{d}\eta_{j_{a}}=-\sum_{b=1}^{s}\mathrm{d}y_{j_{b}}\wedge\sum_{a=b}^{s}\mathrm{d}\eta_{j_{a}}.$
@@
Thus, the coordinates $(y_{j_{a}},\theta_{j_{a}})$ are Darboux coordinates on $S(\mathcal{O})_{\{j\}}$. The coordinates $y_{j_{a}}$ are also coordinates on the stratum $B(\mathcal{O})_{\{j\}}$ of $B(\mathcal{O})$. Complementary coordinates $\theta_{j_{a}}$ are coordinates on the tori $\pi^{-1}(y)$ and are angle coordinates. Now let us show that the multi-time dynamics is linear in $\theta$.
5. The multi-time dynamics of Calogero–Moser–Sutherland system
5.1. The multi-time dynamics on $T^{*}G$
Let $F$ be a polynomial $G$-invariant function on $\mathfrak{g}^{*}$ and $v_{F}$ be the Hamiltonian vector field generated by the pull-back of $F$ to $T^{*}G\simeq\mathfrak{g}\times G$ with respect to the projection to the first factor. For $G=SU(n)$
(95) $v_{F}(x,g)=-i\sum_{k,j+1}^{n}(\nabla Fg)_{kj}\frac{\partial}{\partial g_{kj}}.$
Here $\nabla F(x)$ is the gradient of $F$. It is an $su(n)$-valued function on $su(n)^{*}$, defined as $(y,\nabla F(x))=\frac{d}{dt}F(x+ty)|_{t=0}$, where $(y,x)={\rm tr}(xy)$. The flow line of this vector field, passing theorugh $(x,g)$ at $t=0$ is
(96) $(x(t),g(t))=(x,e^{t\nabla F(x)}g)\ \ .$
Any two such functions $F_{1},F_{2}$ Poisson commute with respect to the standard Poisson bracket on $T^{*}G$, $\{F_{1},F_{2}\}=0$. As a consequence, the corresponding vector field commute $v_{F_{1}},v_{F_{2}}]=0$ and thus, for any $k$ such functions $F_{1},\dots,F_{k}$ the multi-time flow is is defined with flow lines
(97) $(x(\mathbf{t}),g(\mathbf{t}))=(x,e^{i\sum_{j=1}^{k}t_{j}\nabla F_{j}(x)}g)\ \ .$
This dynamics preserves the pull-back to $T^{*}G$ of any $G$-invariant function $H$ on $\mathfrak{g}^{*}$ and the moment map $\mu(x({\bf t}),g({\bf t}))=\mu(x,g)=x-g^{-1}xg$.
5.2. The multi-time dynamics on $S(\mathcal{O})_{\{j\}}$
As we proved in Corollary 2 on the stratum $S(\mathcal{O})_{\{j\}}$, where $\{j\}=\{j_{0}=1<j_{1}<\dots<j_{s}\leq n\}$, the matrix $g$ can be written as
(98) $g=\alpha PY.$
In coordinates $(y,\theta)$, described earlier, the restriction (the pull-back) of the symplectic form $\Omega$ to this stratum is
(99) $\Omega=-\sum_{a=1}^{s}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\theta_{j_{a}},$
which means that Poisson brackets between coordinate functions are:
(100) $\{y_{j_{a}},y_{j_{b}}\}=0,\qquad\{\theta_{j_{a}},\theta_{j_{b}}\}=0,\qquad\{\theta_{j_{a}},y_{j_{b}}\}=\delta_{ab}.$
For a function $F$ which is the pull-back to $T^{*}G$ of a $G$-invariant function on $\mathfrak{g}^{*}$, the Hamiltonian vector fields $v_{F}$ being restricted to $S(\mathcal{O})_{\{j\}}$ can be written as
(101) $v_{F}^{\{j\}}=\sum_{i=1}^{n}\sum_{a=1}^{s}\frac{\partial F}{\partial x_{i}}\frac{\partial x_{i}}{\partial y_{j_{a}}}\frac{\partial}{\partial\theta_{j_{a}}},\qquad\text{ for }k=2,\ldots,n.$
Using (77) we can express $x_{i}$ in terms of $y_{i}$. This gives
(102) $v_{F}^{\{j\}}=\sum_{a=2}^{s}u^{F,\{j\}}_{a}(y)\frac{\partial}{\partial\theta_{j_{a}}},$
@@ 5.2. The multi-time dynamics on $S(\mathcal{O})_{\{j\}}$
(103) $u^{F,\{j\}}_{a}(y)=\frac{n+1-j_{a}}{n}\sum_{i=1}^{j_{a}-1}\frac{\partial F}{\partial x_{i}}(x(y))-\sum_{i=j_{a}}^{n}\frac{\partial F}{\partial x_{i}}(x(y)).$
Here the function $x(y)$ is given by (79).
As a consequence, the multi-time dynamics in $(y,\theta)$ coordinates is
(104) $y_{j_{a}}(\mathbf{t})=y_{j_{a}},\ \ \theta_{j_{a}}(\mathbf{t})=\theta_{j_{a}}-\sum_{k=1}^{s}u^{F_{k},\{j\}}_{a}(y)t_{k},$
where $\mathbf{t}=(t_{1},\dots,t_{s})$. The dynamics is linear in angle variables, as it should be.
Appendix A Lemma about the determinant of the sum of a matrix plus a rank $1$ matrix
@@
Let $D$ be a diagonal matrix $D_{ab}=D_{a}\delta_{ab}$ and $P_{ab}=\varphi_{a}\psi_{b}$ be a 11–dimensional projector. Then
(105) $\det(D-z-P)=\prod_{a}(D_{a}-z)-\sum_{b}\varphi_{b}\psi_{b}\prod_{\begin{subarray}{c}a\neq b\end{subarray}}(D_{a}-z).$
@@
(108) $=\prod_{\alpha}(D_{\alpha}-z)\exp\left(\sum_{k\geq 1}\frac{1}{k}\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}\right).$
@@
(110) $=\sum_{a_{1},\ldots,a_{k}}(D_{a_{1}}-z)^{-1}\cdots(D_{a_{k}}-z)^{-1}\prod_{i=1}^{k}\varphi_{a_{i}}\psi_{a_{i}}.$
@@
(111) $\exp\left(\sum_{k=1}^{\infty}\frac{1}{k}\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}\right)=1-\sum_{b=1}^{n}(D_{b}-z)^{-1}\varphi_{b}\psi_{b}.$